OpenStudy (anonymous):

A car is approaching a traffic light at 18.0 m/s when the light turns red. The driver takes 0.80 s to react and puts his foot on the brake pedal after which the car decelerates at 2.0 m/s^2 . What is the car's stopping distance?

OpenStudy (bradely):

Use formula s= ut+(12)at^2 u=18 m/s t =0.8s a= 2 m/s^2 Source: http://www.mathskey.com/question2answer/

OpenStudy (yttrium):

Total stopping distance may be represented by equation s = s1 + s2 where s1 represents the distance traveled by the car before the he reacts. This is also represented by equation s1 = vt = 18 (0.8) Also, s2 represents the distance traveled by the car when he reacted. You may use the kinematic equation Vf = Vo + 2as, And hence, s2 = (Vf - Vo)/2a where, Vf = 0 (because the car stops), Vo = 18m/s and a = - 2m/s^2 (negative since it is decelerating) Substituting the values, you'll have: s2 = (0-18)/2(-2) Just add s1 and s2 to solve for the total stopping distance s.